OpenStudy (anonymous):

if p-q/p = 2/7 then q/p = ? a 2/5 b 5/7 c 1 d 7/5 e 7/2

OpenStudy (anonymous):

Yes, and the answer is 5/7..... when i do the cross multiplication i get 7/5.

OpenStudy (asnaseer):

sorry - misread your question, should be:\[\frac{p-q}{p}=\frac{p}{p}-\frac{q}{p}=1-\frac{q}{p}\]

OpenStudy (asnaseer):

which does indeed lead to the solution 5/7

OpenStudy (asnaseer):

can you see how?

OpenStudy (anonymous):

NO I'm still confused :-/

OpenStudy (asnaseer):

ok, I've shown you that:\[\frac{p-q}{p}=1-\frac{q}{p}\]so now we can write:\[1-\frac{q}{p}=\frac{2}{7}\]therefore:\[1-\frac{2}{7}=\frac{q}{p}\]\[\frac{7-2}{7}=\frac{q}{p}\]\[\frac{5}{7}=\frac{q}{p}\]

OpenStudy (asnaseer):

make sense?

OpenStudy (anonymous):

Thanks that does make sense!

OpenStudy (asnaseer):

yw